OpenStudy (help108):

1 math quesiton MEDAL AND FAN!!!!!!!!! Consider the following equations: −x − y = 1 y = x + 3 If the two equations are graphed, at what point do the lines representing the two equations intersect? (−1, 2) (−2, 1) (1, −2) (2, −1)

OpenStudy (gniiij):

just graph them go onto mathway.com

OpenStudy (gniiij):

if you need help graphing the first one, you isolate the y

OpenStudy (help108):

@AloneS what do u think?

rishavraj (rishavraj):

see the point of intersection would be common in both lines

OpenStudy (jjcoolja12):

Problem: Solve −x−y=1;y=x+3 Steps: I will try to solve your system of equations. y=x+3;−x−y=1 Step: Solvey=x+3for y: Step: Substitutex+3foryin−x−y=1: −x−y=1 −x−(x+3)=1 −2x−3=1(Simplify both sides of the equation) −2x−3+3=1+3(Add 3 to both sides) −2x=4 −2x −2 = 4 −2 (Divide both sides by -2) x=−2 Step: Substitute−2forxiny=x+3: y=x+3 y=−2+3 y=1(Simplify both sides of the equation) Answer: x=−2 and y=1

rishavraj (rishavraj):

so jst substitute the valus of y from equation 1 in 2 and get x

OpenStudy (jjcoolja12):

B

OpenStudy (help108):

@rishavraj do u think its b?

OpenStudy (-eragon-):

If you change -x-y=1 into workable y=mx+b, you should get y=-x-1, and then you can graph both equations

OpenStudy (-eragon-):

I believe it is b